A 12-inch-thick plain concrete wall has no openings and measures 20 feet by 8 feet. The temperature inside the wall is 68 degrees Fahrenheit and the temperature outside is 7 degrees Fahrenheit. How many square feet of hot-water radiation is required to supply heat equal to the loss?

A. 31.08 B. 27.82 C. 16.59 D. 14.96
Answer #1

rate of heat transfer H = dQ/dt = KA(T2-T1)/L

where K = coeffficient of thermal conductivity

A = area

T2-T1 = change of temperartue

K here = 0.8

so

area needed = 20*8 = 160 ft^2